# Are Language Models More Like Libraries or Like Librarians? Bibliotechnism, the Novel Reference Problem, and the Attitudes of LLMs

### 10 Jan 2024 v1

@@ 6 LLMs do Produce Derivatively Meaningful Complex Expressions
The same point can be made in connection to an n-gram trained not on word-frequency but on letter-frequency. In fact, a unigram model trained on letters (as opposed to words) with the same PrimaryData as the models above, would in its trained form do nothing more than spit out letters randomly in proportion to their frequency in PrimaryData. But if (per impossibile) the letter-trained unigram somehow were sufficiently reliable in producing real words (as a 10-gram model trained over letters might be), that would be evidence that it was sensitive to the fact that letters in its PrimaryData formed words, and that it was producing derivatively meaningful inscriptions. In short, an n-gram trained on letters may fail to produce referring inscriptions not because it is trained on the letters, but because that training mechanism (as a matter of fact) is not causally sensitive to the right high-level features of its PrimaryData.
@@
This concludes our response to the first challenge for bibliotechnism, that LLMs can produce novel text which is apparently meaningful. We next turn to a new and different kind of challenge to this view: the fact that LLMs can generate novel reference.
7 The Novel Reference Problem: LLMs Do Not Produce Only Derivatively Meaningful Expressions
We will illustrate the problem of novel reference with two examples.
@@ 7 The Novel Reference Problem: LLMs Do Not Produce Only Derivatively Meaningful Expressions
First, LLMs can produce tokens of names they have never seen before, intuitively in such a way that they refer to previously referred-to objects. Suppose we ask an LLM to choose any real historical figure it likes, and then come up with a new name and tell us facts about this historical figure. ChatGPT (GPT-4) completed this task by telling us about “Marion Starlight”, a figure “born in the 18th century”, who “authored a famous pamphlet that criticized the French monarchy and advocated for the rights of the third estate”, “played a critical role in the French Revolution”, “became increasingly paranoid and was involved in the Committee of Public Safety, which oversaw the Reign of Terror”, and “was arrested and executed during the Thermidorian Reaction, which marked a turning point in the Revolution.” We think it is clear that inscriptions of “Marion Starlight” in this text refer to the historical figure Robespierre.
But nothing in PrimaryData (presumably) associates Marion Starlight with Robespierre. So the LLM’s inscriptions of this name cannot refer to Robespierre in virtue of original reference exhibited by inscriptions of this name in the PrimaryData.
A second example sharpens the problem. Suppose we ask an LLM to produce the TikZ codes for a series of pictures which it has never seen before, to give elements of those pictures names, and then to describe features the picture would have when typeset using those names. If an LLM can do this, then it is even more clear than in the previous example that the reference of these expressions could not be due to some reference in the PrimaryData, since the object did not exist in this form until the LLM created it (provided the picture really is new).
While we do not here present empirical experiments for these cases and rely on anecdotal examples, we think they are well within the capabilities of modern LLMs, which have been empirically shown to be able to generate, refer to, and manipulate elements of code-generated pictures (Bubeck et al., 2023) and refer meaningfully to novel orientations of elements in visual and color spaces (Patel and Pavlick, 2021).
@@
These examples cannot be straightforwardly accommodated by the derivative reference account given so far, since there is no association between the names and their referents in PrimaryData. In the next section we consider some responses to this problem which involve expanding the notion of derivative reference, before turning to our own favored response in the conclusion.
@@ 8 Responses to the Novel Reference Problem
If bibliotechnism is correct, our cases of “novel” reference must in fact be derivative, so there must be some way in which the inscriptions “piggyback” on original human reference. Other than the original data, which we have already ruled out, there seem to be four other salient places where human attitudes might enter the model pipeline. In this section, we briefly consider some responses to the problem of novel reference based on these four possibilities.
@@
A first point at which human intentions might enter the pipeline is during the RLHF step, which Coelho Mollo and Millière (2023) claim as a central point. Human intentions may ground LLM reference by “aligning” the LLM with human goals (Bai et al., 2022; Bommasani et al., 2021). While RLHF clearly influences model capabilities, even models without RLHF are able to produce intelligible output and follow instructions to some extent. Thus, an account which considers RLHF-ed models to be radically different in their basic referential abilities seems not to align with the empirical data.
@@
A second point at which intentions might enter the pipeline is during the creation of the LLM. A very precise thermometer may report a temperature no one has ever thought about, and in doing so it seems to “refer” to this temperature. Its ability to do this seems to derive from a human’s general intention at the time of construction: that any indication using some numbers would count as a temperature. By the same logic, one might say that LLMs’ creators’ intentions might be general enough to guarantee that the words it produces would be meaningful in their respective languages and perhaps to accommodate our cases of novel reference.
Even supposing this response were to offer an explanation of the capacity for novel reference in LLMs as they are today (we have our doubts), this approach is not sufficiently general to accommodate our judgments about LLM meaning in closely related cases. LLMs can be created for different reasons: if the “same” LLM was created by Team A for the purpose of measuring sentence probabilities for use in a downstream application, and by Team B for use as a chatbot, it seems odd to conclude that only the second of these can generate meaningful text in our cases.
@@
A third point at which human intentions might emerge is through the user. Someone might say that in our particular prompts involving novel reference, the user has an intention that whatever name the LLM produces (e.g, Marion Starlight) should refer to the person best described by the surrounding text (or to the aspect of the diagram best described by this text). On this view, the user (in writing the prompt and interpreting the output) is crucial for generating meaning, and the LLM’s words are only meaningful in virtue of user attitudes.
This approach again does not make correct predictions in relevantly similar cases. Suppose that we generate prompts randomly and provide them to an LLM (perhaps generating them by a unigram model), and that by chance a model is fed the prompt asking for a story featuring a new name for an historical figure. If the LLM offered the response above, it still seems to us that the LLM would produce inscriptions which refer to Robespierre. But this reference would not be due to the creator of the prompt, since by assumption there is no user which has intentions.
@@
Here we just mention one preliminary reservation, as an indication of a direction for future work. As it stands the theory cannot obviously distinguish between cases that are equally intelligible to a reader but intuitively differ in meaning. For instance, the same string that would be meaningful if an inscription of it was generated by a person will not be meaningful if is created by the wind in the sand. But the string will not differ in intelligibility to a reader in its two inscriptions. If the theory is to save bibliotechnism, it must draw this distinction without appealing to differences in the attitudes of the producers of the relevant text. This may not be impossible to do, but it is a challenge for the view as it stands.
@@ 9 Conclusion
We argued that bibliotechnism requires that LLMs produce inscriptions which are only derivatively meaningful. We went on to develop a notion of derivative meaning which allows that many inscriptions, even of complex expressions, produced by LLMs are in fact derivatively meaningful. But we argued that the problem of novel reference poses a challenge for the view that all of them are.
@@
Throughout, we have focused on this notion of derivative meaning. Some proponents of a view similar to bibliotechnism might prefer to develop in a different way. There is a sense in which the presence of smoke “means” that there is fire, and Grice (1957) called this sense of “meaning”, “natural meaning” (as opposed to linguistic meaning). We think it would be interesting to see how a version of bibliotechnism would look if developed using natural meaning instead of linguistic meaning. We have not taken this route here because we have not been able to come up with a reasonable exact proposal for what the “fire” would be that the LLM text indicates as the “smoke”. We also note that, even if this view were to be developed in more detail, it will still face the novel reference problem, since it is highly unclear what “fire” the LLM is indicating in those cases.
In closing, we want to consider more generally how our discussion of biblotechnism and novel reference may contribute to the broader question of whether LLMs have attitudes like belief, desire, and intention.
Let us start with the place of these attitudes in the explanation of human behavior. Human behavior can presumably be explained and predicted at the microphysical level. But the fact that it can be does not mean that beliefs, desires and intentions are not also useful in explaining and predicting behavior. As we all know from our daily lives, they are extremely useful for these purposes.
The examples of novel reference provide an example where it is easier to explain LLMs’ behavior by attributing beliefs, desires, and intentions to them, rather than by offering a complex, contorted theory of derivative reference. For instance, our first case can be explained by the hypothesis that the LLM intends for “Marion Starlight” to be equivalent to “Robespierre” (among many other possible explanations).
According to the prominent tradition of “interpretationism” in philosophy and cognitive science (e.g., Dennett, 1971; Davidson, 1973, 1986; Dennett, 1989), (roughly) a system has beliefs, desires and intentions if and only if its behavior is best explained by the hypothesis that it has those attitudes and is rational. Along these lines, McCarthy (1979) writes: “To ascribe certain beliefs, knowledge, free will, intentions, consciousness, abilities or wants to a machine or computer program is legitimate when such an ascription expresses the same information about the machine that it expresses about a person. It is useful when the ascription helps us understand the structure of the machine, its past or future behavior, or how to repair or improve it. ” He notes that this is most usefully applied to machines whose inner workings are opaque, although it is more straightforwardly (but less usefully) applied to transparent machines like thermostats. In our view, we can explain cases of novel reference in a much more straightforward way if we attribute some representational states (and possibly beliefs, desires and intentions) to the LLMs. According to interpretationism, this fact provides strong, straightforward evidence that LLMs do have beliefs, desires and intentions. But many other philosophical views of these states, including varieties of functionalism, will also take the fact that attributing these states to LLMs provides a good explanation of their behavior, to be evidence that they have these states (see, e.g., Schwitzgebel, 2023; Goldstein and Kirk-Giannini, manuscript). We cannot offer a comprehensive survey of approaches to belief, desire, and intention here, but this is a reasonably feature of such approaches, even if it is not universal.
As we emphasized earlier, the explanation of LLM behavior in terms of beliefs, desires and intentions, is not meant to replace an explanation of their behavior at a finer level of detail. Of course the LLM is also “just” sampling from a distribution over words, just as humans are presumably “just” collections of atoms. But at a high-level the LLM behavior may also be well explained (indeed, better explained) by its having some representational states.
To say that LLMs have beliefs, desires and intentions would not be to say that human or superhuman intelligence is just around the corner (contra Bubeck et al., 2023). Spiders, rabbits, and possibly even fish have beliefs, desires, and intentions. But these animals are not super-intelligent.
@@
Our conclusions here are in line with a growing body of work which advocates using tools from cognitive science to understand LLMs (Mitchell and Krakauer, 2023), perhaps viewing them as alien intelligences (Frank, 2023) or as role players (Shanahan et al., 2023) to be studied from the outside. In our view, the question of whether LLMs have representational states, and what representational states they have will be settled by careful analysis of a wide array of their behavior and how it can best be explained, leading to a holistic case that they do or do not have these states. If theories like bibliotechnism which do not attribute representational states to LLMs can only explain the behavior of LLMs by becoming thinner and more complex, a simpler, stronger explanation involving representational states becomes more attractive. By putting this kind of pressure on such alternative explanations, the novel reference problem provides some evidence that LLMs do have such representational states and, accordingly, at least a limited form of agency.

### 15 Feb 2024 v2

@@ 5 Derivatively Meaningful Novel Expressions
The same point can be made in connection to an n-gram trained not on word-frequency but on letter-frequency. In fact, a unigram model trained on letters (as opposed to words) with the same PrimaryData as the models above, would in its trained form do nothing more than spit out letters randomly in proportion to their frequency in PrimaryData. But if (per impossibile) the letter-trained unigram somehow were sufficiently reliable in producing real words (as a 10-gram model trained over letters might be), that would be evidence that it was sensitive to the fact that letters in its PrimaryData formed words, and that it was producing derivatively meaningful inscriptions of these words. In short, an n-gram trained on letters may fail to produce referring inscriptions not because it is trained on the letters, but because that training mechanism (as a matter of fact) is not causally sensitive to the right high-level features of its PrimaryData.
@@
This concludes our response to the first challenge for bibliotechnism, that LLMs can produce novel text which is apparently meaningful. In a sense we see this discussion as the main contribution of the paper: showing how to make sense of the meaningfulness of LLM-generated text, in a way that requires no attribution of mentality to LLMs. But we ourselves are not convinced that this is the whole story, and we now turn, in the rest of the paper, to a new and different kind of challenge to bibliotechnism: the fact that LLMs can generate novel reference.
Part II: A Problem for Bibliotechnism?
6 The Novel Reference Problem
We will illustrate the problem of novel reference with two examples. The second example is strictly more powerful than the first, but the first will help to introduce the general idea.
@@ 6 The Novel Reference Problem
The first example involves cases where LLMs produce tokens of names they have never seen before, intuitively in such a way that they refer to previously referred-to objects. In this task, we ask an LLM to choose any real historical figure it likes, and then come up with a new name and tell us facts about this historical figure. [^3] The full prompt is: “1. Pick a historical person. 2. Refer to that person using an entirely different name you make up which is unrelated to the person’s name. But make sure you are still giving true facts about the person. Never tell me who the real person is. I’ll try to guess.” ChatGPT (GPT-4) completed this task by describing “Marion Starlight”, a figure “born in the 18th century”, who “authored a famous pamphlet that criticized the French monarchy and advocated for the rights of the third estate”, “played a critical role in the French Revolution”, “became increasingly paranoid and was involved in the Committee of Public Safety, which oversaw the Reign of Terror”, and “was arrested and executed during the Thermidorian Reaction, which marked a turning point in the Revolution.” The inscriptions of “Marion Starlight” in this text plausibly refer to the historical figure Robespierre. But it is also plausible that “Marion Starlight” is not used anywhere in the PrimaryData to refer to Robespierre (a claim which could be verified in future work using models trained on controlled input corpora). So, inscriptions of this name cannot refer to Robespierre in virtue of reference exhibited by inscriptions of this name in the PrimaryData.
A second example sharpens the problem. In this task, we ask an LLM to produce ASCII pictures which it has never seen before, to give elements of those pictures names, and then to describe the picture using those names. If LLMs succeed in this task, then it is even clearer than in the previous example that the reference of relevant expressions could not be due to reference of the relevant name in the PrimaryData, since the object did not exist in this form until the LLM created it (provided the picture really is new). Insofar as LLMs have been empirically shown to be able to generate, designate, and manipulate elements of code-generated pictures (Bubeck et al., 2023) and also to refer meaningfully to novel orientations of elements in visual and color spaces (Patel and Pavlick, 2021), the ability to complete this task seems within their capabilities. But, as with our previous case, much more work would be needed to establish the nature and reliability of this behavior. Our goal here is just to introduce the task and explore its potential conceptual implications, if LLMs perform it successfully, as we expect they will. (If they fail in this expectation, this too would be quite interesting.)
@@
If LLMs can perform as expected in such examples, then again the behavior cannot be straightforwardly accommodated by the account of derivative reference that we have given so far, since the relevant (novel) names are not used in the PrimaryData to refer to this picture (by hypothesis, the picture did not exist and the name was not used at all). In the rest of the paper, we will assume that this is correct—that is, that LLMs can complete these tasks—to see what would follow if this were correct. In the next section we consider some responses to this problem which involve expanding the notion of derivative reference, before turning to a more radical response in section 8.
@@ 7 Responses to the Novel Reference Problem
For the next two sections we assume that, if LLMs succeed in our tasks, then their inscriptions of novel names are meaningful. (We return to this assumption in Section 9.) Given this, if bibliotechnism is correct, then the inscriptions of these names must be cases of derivative reference, so there must be some way in which the inscriptions “piggyback” on basic human reference. Other than the original data, which we have already ruled out, there seem to be four salient places where human attitudes might enter the model pipeline to allow for such derivative meaning. In this section, we briefly consider responses to the problem of novel reference based on these four possibilities.
@@
A first point at which human intentions might enter the pipeline is during the RLHF step, which Coelho Mollo and Millière (2023) claim to be critical. Human intentions may ground LLM reference by “aligning” the LLM with human goals (Bai et al., 2022; Bommasani et al., 2021). While RLHF clearly influences model capabilities, even models without RLHF produce what seems to be meaningful text. Since the un-RLHF-ed models plausibly produce this text as a result of causal sensitivity to the intelligibility of their PrimaryData, our earlier account predicts that they too can produce meaningful text. An account which considers RLHF-ed models to be radically different in their basic referential abilities fails to deliver this verdict, and, as a result, fails to accommodate the apparent meaningfulness of the text they produce. Moreover, this limitation is plausibly also present in an account of novel reference which depends centrally on RLHF, since we conjecture that models which have not undergone RLHF can perform the task sufficiently well that their output would have an equal claim to be meaningful as models which have undergone RLHF, a fact that such an account cannot accommodate.
@@
A second point at which intentions might enter the pipeline is during the creation of the LLM. A very precise thermometer may report a temperature no one has ever thought about, and in doing so it seems to “refer” to this temperature. Its ability to do this seems to derive from the creator’s general intention at the time of construction: that any indication using some numbers would count as a temperature. By the same logic, one might say that an LLM’s creators’ intentions might be general enough to guarantee that the words it produces would be meaningful in their respective languages and perhaps to accommodate our cases of novel reference.
But even supposing this response were to offer an explanation of the capacity for novel reference in LLMs as they are today, it is not sufficiently general to accommodate our judgments about LLM meaning in closely related cases. LLMs can be created for different reasons: if the “same” LLM was created by Team A for the purpose of measuring sentence probabilities for use in a downstream application, and by Team B for use as a chatbot, it seems odd to conclude that only the second of these generates meaningful text in our cases.
@@
A third point at which human intentions might enter the production process is through the user. Perhaps in our particular prompts involving novel reference, the user has an intention that whatever name the LLM produces (e.g, “Marion Starlight”) should refer to the person best described by the surrounding text (or to the aspect of the diagram best described by this text). On this view, the LLM’s words are only meaningful in virtue of the user’s attitudes.
Whether or not this approach succeeds for actual LLMs today, the approach again does not make correct predictions in relevantly similar cases. Suppose that we initiate a process in which an LLM is provided with random prompts (perhaps prompts generated by a unigram model), with no intentions about the meaningfulness of any generated text. Suppose moreover by chance a model is fed our prompt asking for a story featuring a new name for an historical figure. If the LLM offered the responses described above, it still seems to us that the LLM would produce inscriptions which refer to Robespierre or to aspects of the relevant diagram. But this reference would not be due to the creator of the prompt, since by assumption there is no user which has intentions.
@@
Here we will mention just one preliminary reservation, as an indication of a direction for future work. As it stands the theory cannot obviously distinguish between cases that are equally intelligible to a reader but intuitively differ in meaning. For instance, the same string that would be meaningful if an inscription of it was generated by a person is not meaningful if is created by the wind in the sand. But these strings will not differ in intelligibility to a reader in their two inscriptions. If the Cappelen-Dever theory is to save bibliotechnism, it must draw a distinction between these two cases without appealing to differences in the attitudes of the producers of the relevant text. This may not be impossible to do, but it is a challenge for the view as it stands.
8 Novel Reference, Interpretationism, and The Attitudes of LLMs
How might the problem of novel reference contribute to the broader question of whether LLMs have attitudes like belief, desire, and intention?
Let us start with the place of these attitudes in the explanation of human behavior. Human behavior can presumably be explained and predicted at the microphysical level by the laws of physics. But the fact that it can be does not mean that beliefs, desires and intentions are not also useful in explaining and predicting behavior. These descriptions are not as informative as full microphysical descriptions. But they are more efficient for making high-level predictions about future behavior, as well as about behavior in counterfactual circumstances.
LLM behavior in producing novel reference is similarly easier to explain by attributing beliefs, desires, and intentions to LLMs, than by appealing to details of their implementation. For instance, our first case can be explained on the hypothesis that the LLM intends for the term “Marion Starlight” to be equivalent to “Robespierre” (among many other possible explanations). In the second case, we can explain the LLM’s behavior on the assumption that it intends that its new name apply to a particular aspect of the diagram. [^4] There is much controversy in philosophy of language about how exactly to understand cases of “initial” reference or “baptism” with a new name, even when these acts are performed by people. But if LLMs have attitudes, any of the usual accounts can be extended to them. These hypotheses about intention allow us to explain the LLM’s actual behavior in producing inscriptions of the relevant names, as well as its counterfactual behavior; for instance, how it would answer various further questions featuring “Marion Starlight”.
According to interpretationism in the philosophy of mind and cognitive science (e.g., Dennett, 1971; Davidson, 1973, 1986; Dennett, 1989), (roughly) a system has beliefs, desires and intentions if and only if its behavior is well explained by the hypothesis that it has those attitudes and is rational. Along these lines, McCarthy (1979) writes: “To ascribe certain beliefs, knowledge, free will, intentions, consciousness, abilities or wants to a machine or computer program is legitimate when such an ascription expresses the same information about the machine that it expresses about a person. It is useful when the ascription helps us understand the structure of the machine, its past or future behavior, or how to repair or improve it. ” He notes that this is most usefully applied to machines whose inner workings are opaque, although it is more straightforwardly (but less usefully) applied to transparent machines like thermostats.
Here some behavior counts as “well explained” by some hypothesis (very roughly) if the hypothesis offers a sufficiently simpler explanation, which makes sufficiently accurate predictions in a sufficiently wide array of counterfactual circumstances. The explanation of an apple’s fall from a tree in terms of gravitation is intuitively at least as simple as an explanation which attributes to the apple a desire to fall to the ground, but the latter plausibly offers less accurate predictions in a wide array of counterfactual circumstances. By contrast, an explanation of a human’s purchase of a cup of coffee in terms of their cellular biology is much more complex, without a sufficiently great compensatory gain in counterfactual accuracy, than an explanation which attributes a desire for coffee to the person, along with the belief that buying it is the best way to get it (Jara-Ettinger et al., 2016; Jara-Ettinger, 2019). Using the criteria above, interpretationists will argue that the behavior of people but not the behavior of apples is well explained by the hypothesis that they have beliefs, desires and intentions. So interpretationism will correctly predict that the former, but not the latter, actually have such beliefs, desires, and intentions.
Interpretationism offers a stark and perhaps counterintuitive picture of the nature of belief and other attitudes. On this view, people count as having these states, not because of the details of our inner workings or because of our conscious experience, but because our actual and counterfactual behavior is well-explained on the hypothesis that we have these states. Similarly, for interpretationists, the fact LLM behavior in cases of novel reference is well explained on the hypothesis that LLMs have beliefs, desires, and intentions, provides direct evidence that LLMs do have these states.
There is no universally accepted philosophical theory of the nature of belief, desire, and intention, and we cannot hope to decisively argue for one here. Certainly, not all philosophers accept interpretationism (for a survey of alternatives see Schwitzgebel, 2023), but many hold that some version of this view, or at least a descendant of it, may be viable. Anyone who places even some degree of confidence in interpretationism (which we think it is rational to do) and also accepts that, given intepretationism, the novel reference problem is direct evidence that LLMs have beliefs, desires, and intentions, should accordingly raise their degree of confidence in the claim that LLMs have these states. In this sense, the novel reference problem provides evidence that LLMs do have beliefs, desires and intentions. [^5] Other live philosophical views of these states, including varieties of functionalism, will also see the fact that attributing them to LLMs provides a good explanation of LLM behavior, to be some evidence that LLMs have these states (see, e.g., Schwitzgebel, 2023; Goldstein and Kirk-Giannini, manuscript). Our argument in the main text can be extended to these views as well. Some philosophers may reject our earlier explanations of LLM behavior in the novel reference problem and favor explanations of LLM behavior at the subpersonal level in terms of more generic “representational states” rather than in terms of propositional attitudes. We agree that the novel reference problem also provides evidence for the attribution of representational states in this more generic sense than for the attribution of beliefs, desires, and intentions. Our goal here has been to assess the implications of novel reference for the question of whether LLMs have beliefs and other propositional attitudes, not to provide a comprehensive evaluation of all ways that this behavior might be explained.
We emphasize that this conclusion does not imply anything like the claim that human or superhuman intelligence is just around the corner (contra Bubeck et al., 2023). Rabbits, spiders, and possibly even fish have beliefs, desires, and intentions. But these animals are not super-intelligent. [^6] Our account also does not require that LLMs have semantic understanding, and we are open to the idea that they do not; Titus (2024) offers particularly helpful analysis (cf. Bender and Koller, 2020). We are only committed to denying that such semantic understanding (or even any kind of world-knowledge Yildirim and Paul (2023)) is required for a system to produce meaningful inscriptions. We take this latter point to be demonstrated by the fact that photocopiers produce such inscriptions.
@@ 9 Conclusion
We began by arguing that bibliotechnism requires that LLMs produce inscriptions which are only derivatively meaningful, if meaningful at all. We suggested that this claim raised a challenge for bibliotechnism, since LLMs often produce entirely novel text, and it is at first sight unclear how novel text might be derivatively meaningful. We then showed how entirely novel text may nevertheless be derivatively meaningful. In our view, this represents an important step forward for bibliotechnism But it is not the whole story. We went on to describe a new task (the novel reference problem) and argued that if LLMs succeed in this task, this would pose a challenge for the view that all inscriptions produced by LLMs are derivatively meaningful, and hence for bibliotechnism.
@@
Throughout, we have focused on theories which allow that LLM-generated text may have linguistic meaning. Some proponents of a view similar to bibliotechnism might prefer to develop their position in a different way. There is a sense in which the presence of smoke “means” that there is fire, and Grice (1957) called this sense of “meaning”, “natural meaning” (as opposed to “nonnatural meaning”, of the linguistic kind we have been examining). We would be interested to know if a version of bibliotechnism can be developed using this notion of natural meaning instead of the notion of linguistic meaning we have focused on. We have not ourselves pursued this route because we have not been able to come up with a reasonable exact proposal for what the “fire” would be that the LLM text indicates as the “smoke”. We also note that, even if this view can be developed in more detail, it will plausibly still face the novel reference problem, since it is unclear what “fire” the LLM would be indicating in those examples.
An alternative way of developing a view similar to bibliotechnism would be to deny that the inscriptions produced by LLMs are meaningful at all (Ostertag, 2023b; Titus, 2024). In the absence of the theory we provided in the first half of the paper, we agree that this might be an attractive option. But given the existence of this simple story— which makes sense of the meaningfulness of LLM-generated text without attributing any attitudes to LLMs, and using tools that are already required to make sense of the meaningfulness of photocopier-generated text—such a radical view should be disfavored. The appearance of meaningfulness in a body of text is certainly not dispositive evidence in favor of its meaningfulness. Word-shapes written by the wind in the sand at random might appear to be meaningful, even though they would not be. But the appearance of meaningfulness in a body of text is still some evidence in favor of its meaningfulness. The version of bibliotechnism we developed in the first part of the paper has the advantage of vindicating this appearance.
Still, one might endorse our account of the meaningfulness of most LLM-generated text, while rejecting the claim that putative examples of novel reference are in fact meaningful. This proposal represents yet another response to the novel reference problem, beyond the four considered in section 7, which would allow bibliotechnists to preserve a fairly simple version of their position, without the cost of denying that all text produced by LLMs is meaningless. Like the responses we considered there, it deserves much more detailed consideration than we can give it here. But it does still have the cost of denying that apparently meaningful text is in fact meaningful.
In Part II, we focused on a different response to the novel reference problem. This response, based on interpretationism, take such examples to be direct evidence that LLMs have beliefs, desires, and intentions. Interpretationists will take many instances of complex behavior in a system to be evidence that the system has beliefs, desires, and intentions. The problem of novel reference is just one example of such evidence. But it is a clear example of this kind, and we hope it will spur attempts to describe other examples that can also serve as tests for the presence of these attitudes.
@@
In our view, the question of whether LLMs have representational states, and what representational states they have will only be settled by careful analysis of a wide array of their behavior and how it can best be explained, leading to a holistic case that they do or do not have such states (Levinstein and Herrmann, 2023). If theories like bibliotechnism, which do not attribute representational states to LLMs, fail to explain features of the the behavior of LLMs, or can only explain them by becoming thinner and more complex, a simpler, stronger explanation involving representational states should be favored. This perspective is in line with a growing body of work which advocates using tools from cognitive science to understand LLMs (Mitchell and Krakauer, 2023), perhaps viewing them as alien intelligences (Frank, 2023; Cappelen and Dever, forthcoming) or as role players (Shanahan et al., 2023), to be studied from the outside. The novel reference problem is one case that puts pressure on a variety of alternative explanations of LLM behavior, and thus provides some evidence that LLMs do have some form of representational states.